OpenStudy (anonymous):

find the sum of 2sqrt6 and sqrt24

OpenStudy (callisto):

\[2\sqrt6 + \sqrt{24} =2\sqrt6 + \sqrt{2\times2\times6} = 2\sqrt6 + 2\sqrt{6} = 4\sqrt6\]

OpenStudy (kamille):